OpenStudy (anonymous):

How do you find slope when given 5y-8x=7

OpenStudy (anonymous):

rewrite the function till you get y=ax+b; a = slope

OpenStudy (anonymous):

Set the equation to \[y=8/5x+7/5\] (slope intercept form). The slope is 8/5.

hero (hero):

y = (7+8x)/5

OpenStudy (amistre64):

or, when given ax+by=c; slope = -a/b

OpenStudy (amistre64):

--8/5 = 8/5

OpenStudy (anonymous):

So how would you get from point a to b. I know how to find slope intercept for for both x,y and how to find the slope in via slope formula. How would you break it down to find slope in the current form?